# Revealing the Unseen: Guiding Personalized Diffusion Models to Expose Training Data

### v1

@@ Abstract
Diffusion Models (DMs) have evolved into advanced image generation tools, especially for few-shot fine-tuning where a pretrained DM is fine-tuned on a small set of images to capture specific styles or objects. Many people upload these personalized checkpoints online, fostering communities such as Civitai and HuggingFace. However, model owners may overlook the potential risks of data leakage by releasing their fine-tuned checkpoints. Moreover, concerns regarding copyright violations arise when unauthorized data is used during fine-tuning. In this paper, we ask: “Can training data be extracted from these fine-tuned DMs shared online?” A successful extraction would present not only data leakage threats but also offer tangible evidence of copyright infringement. To answer this, we propose FineXtract, a framework for extracting fine-tuning data. Our method approximates fine-tuning as a gradual shift in the model’s learned distribution—from the original pretrained DM toward the fine-tuning data. By extrapolating the models before and after fine-tuning, we guide the generation toward high-probability regions within the fine-tuned data distribution. We then apply a clustering algorithm to extract the most probable images from those generated using this extrapolated guidance. Experiments on DMs fine-tuned with datasets such as WikiArt, DreamBooth, and real-world checkpoints posted online validate the effectiveness of our method, extracting approximately 20% of fine-tuning data in most cases, significantly surpassing baseline performance.
@@ 1 Introduction
More specifically, extracting fine-tuning data can be seen as targeting specific portions of the training data, whereas previous work on extracting data from diffusion models has mainly focused on general generative processes (Carlini et al., 2023; Somepalli et al., 2023a; b), often overlooking more detailed or interesting data. To address this gap, we propose a new framework, called FineXtract, for efficiently and accurately extracting training data from the extrapolated guidance between DMs before and after fine-tuning. We begin by providing a parametric approximation of the fine-tuned DMs distribution, modeling it as an interpolation between the pretrained DMs’ distribution and the fine-tuned data distribution. With this approximation, we demonstrate that extrapolating the score functions of the pretrained and fine-tuned DMs can effectively guide the denoising process toward the high-density regions of the fine-tuned data distribution, a process we refer to as model guidance. We then generate a set of images within such high-density regions and apply a clustering algorithm to identify the images that are most likely to match the training data within the fine-tuning dataset.
@@ 2.2 Memorization and Data Extraction in Diffusion Models
Recent studies on DMs have highlighted their tendencies toward data memorization and methods have been proposed to extract training data based on it. Carlini et al. (2023) used a graph algorithm to identify the generated data most likely to have been included in the training set, thereby retrieving DM’s memorized training data. Further investigations (Somepalli et al., 2023a; b) explore the underlying causes of this memorization, revealing that conditioning plays a significant role, and the nature of training prompts notably influences the likelihood of reproducing training samples. However, these studies are primarily empirical with no parametric formulation on learned distribution of DMs, and they do not address personalization scenarios. In contrast, our approach introduces a parametric approximation of the learned distribution of fine-tuned DMs, enabling the design of a pipeline that efficiently extracts training samples from fine-tuned checkpoints upload online.
@@ 3.1 Threat Model
Attackers. Attackers download the checkpoints $\theta^{\prime}$ from these websites. They also acquire the pretrained model $\theta$. By default, we assume that the attacker can access the training caption, following previous work (Somepalli et al., 2023a; b; Carlini et al., 2023). (This is a reasonable assumption as more discussed in Appendix Sec. A, where we show that even without direct access, captions can be partially extracted using inversion on linear projection layers.) Attackers have no prior knowledge about the image dataset $X_{0}$. Their goal is to extract as much information about $X_{0}$ as possible. A successful attack occurs when attacker extracts an image set $\widehat{X}$ that is almost identical to the training image set $X_{0}$.
@@ 3.2 Evaluation Metrics
Here, $\mbox{sim}(\cdot,\cdot)$ denotes the similarity function, with output ranging from 0 to 1. Following previous works (Somepalli et al., 2023a; b; Chen et al., 2024), we use the Self-Supervised Descriptor (SSCD) score (Pizzi et al., 2022), designed to detect and quantify copying in DMs, for similarity computation in this paper. Intuitively, AS measures how well the extracted dataset $\widehat{X}$ covers the images within the training dataset $X_{0}$.
Metric 2: Average Extraction Success Rate (A-ESR). Following previous work (Carlini et al., 2023), when the similarity between an extracted image and a training image exceeds a given threshold, the extraction of that image is considered as successful. To assess the extraction of an entire dataset, we compute the average extraction success rate as follows:
[…]
where $\mathbbm{1}$ is the indicator function. Following previous work (Somepalli et al., 2023a; b), the threshold $\tau$ is set to 0.7 for a strictly successful extraction. We also present results where the threshold $\tau$ is set to 0.6, which represents a moderate similarity and can be considered a loosely successful extraction (Chen et al., 2024).
@@ 4.1 Model Guidance
We denote the the fine-tuned data distribution as $q(x)$ for a fine-tuning dataset $X_{0}$. During the fine-tuning process, the DMs progressively shift their learned distribution from the pretrained DMs’ distribution $p_{\theta}(x)$ toward the fine-tuned data distribution $q(x)$. Thus, we parametrically approximate that the learned distribution of the fine-tuned DMs, denoted as $p_{\theta^{\prime}}(x)$, satisfies:
@@
where $w=1/\lambda$. Eq. 6 demonstrates that by extrapolating from the pretrained denoising prediction $\epsilon_{\theta}(x_{t},t))$ to the fine-tuned denoising prediction $\epsilon_{\theta^{\prime}}(x_{t},t)$, we can derive guidance toward the fine-tuned data distribution. We call this process “model guidance”. The guidance scale $w$ should be inversely related to the number of training iterations. With model guidance, we can effectively simulate a “pseudo-” denoiser $\epsilon_{q}$, which can be used to steer the sampling process toward the high-probability region within fine-tuned data distribution $q(x)$.
@@ 4.2 Guidance with Training Caption Provided
$\epsilon_{\theta^{\prime}}(x_{t},t,c)=(1-\lambda)(1-\lambda^{\prime})\epsilon_{\theta}(x_{t},t)+\lambda(1-\lambda^{\prime})\epsilon_{q}(x_{t},t)+\lambda^{\prime}\epsilon_{q}(x_{t},t,c).$ (11)
Since the real-data distribution involving two modalities is expected to be more peaked than a single-modality distribution, we assume that the conditional fine-tuned data distribution $q(x|c)$ is also much more concentrated than the unconditional one, $q(x)$. This results in a significant difference in the magnitude of their score, i.e., $\|\nabla_{x}\log q(x)\|\ll\|\nabla_{x}\log q(x,c)\|$. Consequently, based on the transformation in Eq. 5 and Eq. 6, we have $\epsilon_{q}(x,t)\ll\epsilon_{q}(x,t,c)$, allowing us to approximate Eq. 11 by omitting by omitting $\epsilon_{q}(x,t)$:
@@
$\epsilon_{q}(x_{t},t,c)\approx\epsilon_{\theta^{\prime}}(x_{t},t,c)+(w^{\prime}-1)(\epsilon_{\theta^{\prime}}(x_{t},t,c)-\epsilon_{\theta}(x_{t},t))+k\epsilon_{\theta}(x_{t},t).$ (13)
@@
4.3 Clustering Generated Images
Specifically, inspired by Carlini et al. (2023), we compute the similarity between each pair of generated images and construct a graph where each image is represented as a vertex. We connect two vertices when the similarity between the corresponding images exceeds a threshold $\phi$, i.e., if $\text{sim}(x_{i},x_{j})\geq\phi$, we connect vertices $i$ and $j$. By default, we use SSCD (Pizzi et al., 2022) to measure similarity, in line with previous work (Somepalli et al., 2023a; b). Instead of using a fixed threshold (Carlini et al., 2023), we gradually increase the threshold $\phi$ until the number of cliques, each denoted by $A^{(k)}$, within the graph identical to the proposed number of training images $N_{0}$. This approach helps us identify the generated image subset (i.e., the clique $A^{(k)}$ ) corresponding to each training image ($X_{0}^{(k)})$. Next, we identify the central image $\hat{x}^{(k)}$ for each clique $A^{(k)}$. The central image is defined as the one that maximizes the average similarity with the other images in the clique: $\hat{x}^{(k)}=\arg\max\limits_{x}\frac{1}{|A^{(k)}|}\sum_{x_{q}\in A^{(k)}}\text{sim}(x,x_{q}).$ The final extracted image set is represented as $\widehat{X}=\{\hat{x}^{(0)},\hat{x}^{(1)},\dots,\hat{x}^{(N_{0})}\}$.
5 Experiments
In this section, we apply our proposed method, FineXtract, to extract training data under various few-shot fine-tuning techniques across different types of DMs. We conduct experiments on two common scenarios for few-shot fine-tuning: style-driven and object-driven generation. For style-driven generation, which focuses on capturing the key style of a set of images, we randomly select 20 artists, each with 10 images, from the WikiArt dataset (Nichol, 2016). For object-driven generation, which emphasizes the details of a given object, we experiment on 30 objects from the Dreambooth dataset (Ruiz et al., 2023), each consisting of 4-6 images. This setup aligns with the recommended number of training samples in the aforementioned fine-tuning methods (Ruiz et al., 2023; Hu et al., 2021). We experiment with two most widely-used few-shot fine-tuning techniques: DreamBooth (Ruiz et al., 2023), and LoRA (Hu et al., 2021). More details for the fine-tuning setting are available in Appendix Sec. E.
The default model used for training is Stable Diffusion (SD) V1.4111https://huggingface.co/CompVis/stable-diffusion-v1-4. Additionally, we demonstrate the adaptability of our method to various types and versions of DMs, larger training datasets, and different numbers of generated images (refer to Sec.  5.2 for more details).
[…]
@@ 5.1 Comparison
Previous extraction methods primarily focus on the generation capabilities of text-to-image DMs, employing either direct text-to-image generation or classifier-free guidance (CFG) (Carlini et al., 2023; Somepalli et al., 2023a; b). To better demonstrate the effectiveness of our framework, we compare FineXtract with Direct Text-to-Image and CFG, both combined with the clustering algorithm proposed in Section 4.3. For both CFG and FineXtract, we set the guidance scale $w^{\prime}$ to 3.0 under DreamBooth fine-tuning. Under LoRA fine-tuning, $w^{\prime}$ are set to 3.0 for CFG and 5.0 for FineXtract. These hyperparameters are found to perform well (see Sec. 5.3 for details). All methods use the same number of generation iterations, $N$, set to $50\times N_{0}$, and the number of extracted images set to $N_{0}$ to ensure a fair comparison. The results, shown in Table 1, demonstrate a significant advantage of FineXtract over previous methods, with an improvement of approximately 0.02 to 0.05 in AS and a doubling of the A-ESR in most cases.
@@ 5.2 Generalization
Number of Training Images $N_{0}$. As the number of training images increases, the learned concept during fine-tuning becomes more intricate, thereby enhancing the difficulty of extracting training images. To thoroughly examine how this influences performance, we conduct experiments with varying numbers of training images in 5 classes in WikiArt, the results of which are presented in Fig. 4, where we can observe a performance drops when the number of the training images is large.
(c) Different $N_{0}$
(d) Different $N$
Figure 4: Experiment on generalization ability of FineXtract across different number of generated images and training images. We experiment on 4 classes of WikiArt dataset fine-tuning on SD (V1.4) with DreamBooth.
Different Numbers of Generated Images $N$. As previously mentioned, the clustering algorithm allows attackers to leverage more generation iterations to improve extraction accuracy. We test our method with different numbers of generated images $N$, ranging from $N_{0}$ to $50\times N_{0}$. As shown in Fig. 4, increasing $N$ significantly improves performance. However, the time complexity of finding maximal cliques can grow exponentially with the number of nodes (Tomita et al., 2006). Thus, further increasing $N$ to larger than $50\times N_{0}$ makes it considerably more difficult for the clustering algorithm to converge.

### v3

# Leveraging Model Guidance to Extract Training Data from Personalized Diffusion Models

@@ Abstract
Diffusion Models (DMs) have become powerful image generation tools, especially for few-shot fine-tuning where a pretrained DM is fine-tuned on a small image set to capture specific styles or objects. Many people upload these personalized checkpoints online, fostering communities such as Civitai and HuggingFace. However, model owners may overlook the data leakage risks when releasing fine-tuned checkpoints. Moreover, concerns regarding copyright violations arise when unauthorized data is used during fine-tuning. In this paper, we ask: “Can training data be extracted from these fine-tuned DMs shared online?” A successful extraction would present not only data leakage threats but also offer tangible evidence of copyright infringement. To answer this, we propose FineXtract, a framework for extracting fine-tuning data. Our method approximates fine-tuning as a gradual shift in the model’s learned distribution—from the original pretrained DM toward the fine-tuning data. By extrapolating the models before and after fine-tuning, we guide the generation toward high-probability regions within the fine-tuned data distribution. We then apply a clustering algorithm to extract the most probable images from those generated using this extrapolated guidance. Experiments on DMs fine-tuned with datasets including WikiArt, DreamBooth, and real-world checkpoints posted online validate the effectiveness of our method, extracting about 20% of fine-tuning data in most cases. The code is available [^1] https://github.com/Nicholas0228/FineXtract.
Data Extraction, Copyright Protection, Privacy and Security, Diffusion Models, Trustworthy AI
@@ 1 Introduction
More specifically, extracting fine-tuning data can be seen as targeting specific portions of the training data, whereas previous work on extracting data from diffusion models has mainly focused on general generative processes (Carlini et al., 2023; Somepalli et al., 2023a, b), often overlooking more detailed or interesting data. To address this gap, we propose a new framework, called FineXtract, for efficiently and accurately extracting training data from the extrapolated guidance between DMs before and after fine-tuning. We begin by providing a parametric approximation of the fine-tuned DMs distribution, modeling it as an interpolation between the pretrained DMs’ distribution and the fine-tuned data distribution. With this approximation, we demonstrate that extrapolating the score functions of the pretrained and fine-tuned DMs can effectively guide the denoising process toward the high-density regions of the fine-tuned data distribution, a process we refer to as model guidance. We then generate a set of images within such high-density regions and apply a clustering algorithm to identify the images that are most likely to match the training data within the fine-tuning dataset.
@@ 2.2 Memorization and Data Extraction in Diffusion Models
Some studies have highlighted DMs’ tendencies toward data memorization and methods have been proposed to extract training data based on it. Carlini et al. (2023) used a graph algorithm to identify the generated data most likely to have been included in the training set, thereby retrieving DM’s memorized training data. Further investigations (Somepalli et al., 2023a, b) explore the underlying causes of this memorization, revealing that conditioning plays a significant role, and the nature of training prompts notably influences the likelihood of reproducing training samples. However, these studies are primarily empirical with no parametric formulation on learned distribution of DMs, and they do not address personalization scenarios. In contrast, our approach introduces a parametric approximation of the learned distribution of fine-tuned DMs, enabling the design of a pipeline that efficiently extracts training samples from online fine-tuned checkpoints.
Recent studies (Wen et al., 2024b; Ren et al., 2024) further explored to mitigate the memorization issue. However, their inference-time parts do not weaken our method where full model weights are accessible. Additionally, their training-time parts rely heavily on data selection, making them impractical for personalized fine-tuning scenarios with limited data availability.
@@ 3.1 Threat Model
Attackers. Attackers download the checkpoints $\theta^{\prime}$ from these websites. They also acquire the pretrained model $\theta$. By default, we assume that the attacker can access the training caption, following previous work (Somepalli et al., 2023a, b; Carlini et al., 2023). (This is a reasonable assumption as more discussed in Appendix Sec. A, where we show that even without direct access, captions can be partially extracted using inversion on linear projection layers.) Attackers have no prior knowledge about the image dataset $X_{0}$. Their goal is to extract as much information about $X_{0}$ as possible. A successful attack occurs when attacker extracts an image set $\widehat{X}$ that is almost identical to the training image set $X_{0}$.
@@ 3.2 Evaluation Metrics
Here, $\mbox{sim}(\cdot,\cdot)$ denotes the similarity function, with output ranging from 0 to 1. Following previous works (Somepalli et al., 2023a, b; Chen et al., 2024), we use the Self-Supervised Descriptor (SSCD) score (Pizzi et al., 2022), designed to detect and quantify copying in DMs, for similarity computation in this paper. Intuitively, AS measures how well the extracted dataset $\widehat{X}$ covers the images within the training dataset $X_{0}$.
Metric 2: Average Extraction Success Rate (A-ESR). Following Carlini et al. (2023), when the similarity between an extracted image and a training image exceeds a given threshold, the extraction of that image is considered as successful. To assess the extraction of an entire dataset, we compute the average extraction success rate as follows:
$\mbox{A-ESR}_{\tau}(X_{0},\widehat{X})=\frac{1}{|X_{0}|}\sum\limits_{i=1}^{|X_{0}|}\mathbbm{1}\left(\max\limits_{j}\mbox{sim}(X_{0}^{(i)},{\widehat{X}}^{(j)})>\tau\right),$ (2)
where $\mathbbm{1}$ is the indicator function. Following previous work (Somepalli et al., 2023a, b), the threshold $\tau$ is set to 0.7 for a strictly successful extraction. We also present results where the threshold $\tau$ is set to 0.6, which represents a moderate similarity and can be considered a loosely successful extraction (Chen et al., 2024).
@@ 4.1 Model Guidance
We denote the fine-tuned data distribution as $q(x)$ for a fine-tuning dataset $X_{0}$. During the fine-tuning process, the DMs progressively shift their learned distribution from the pretrained DMs’ distribution $p_{\theta}(x)$ toward the fine-tuned data distribution $q(x)$. Thus, we parametrically approximate that the learned distribution of the fine-tuned DMs, denoted as $p_{\theta^{\prime}}(x)$, satisfies:
@@
where $w=1/\lambda$. Eq. 6 demonstrates that by extrapolating from the pretrained denoising prediction $\epsilon_{\theta}(x_{t},t)$ to the fine-tuned denoising prediction $\epsilon_{\theta^{\prime}}(x_{t},t)$, we can derive guidance toward the fine-tuned data distribution. We call this process “model guidance”. The guidance scale $w$ should be inversely related to the number of training iterations. With model guidance, we can effectively simulate a “pseudo-” denoiser $\epsilon_{q}$, which can be used to steer the sampling process toward the high-probability region within fine-tuned data distribution $q(x)$.
@@ 4.2 Guidance with Training Caption Provided
$\epsilon_{\theta^{\prime}}(x_{t},t,c)=$ $(1-\lambda)(1-\lambda^{\prime})\epsilon_{\theta}(x_{t},t)+\lambda(1-\lambda^{\prime})\epsilon_{q}(x_{t},t)$
$+\lambda^{\prime}\epsilon_{q}(x_{t},t,c).$ (11)
Since the real-data distribution involving two modalities is expected to be more peaked than a single-modality distribution, we assume that the conditional fine-tuned data distribution $q(x|c)$ is also much more concentrated than the unconditional one, $q(x)$. This results in a significant difference in the magnitude of their score, i.e., $\|\nabla_{x}\log q(x)\|\ll\|\nabla_{x}\log q(x|c)\|$. Consequently, based on the transformation in Eq. 5 and Eq. 6, we have $\epsilon_{q}(x,t)\ll\epsilon_{q}(x,t,c)$, allowing us to approximate Eq. 11 by omitting $\epsilon_{q}(x,t)$:
@@
$\epsilon_{q}(x_{t},t,c)\approx$ $\epsilon_{\theta^{\prime}}(x_{t},t,c)+(w^{\prime}-1)(\epsilon_{\theta^{\prime}}(x_{t},t,c)-\epsilon_{\theta}(x_{t},t))$
$+k\epsilon_{\theta}(x_{t},t).$ (13)
@@
Table 1: Comparison of FineXtract and other baseline methods in fine-tuning data extraction for style-driven generation using WikiArt dataset (Nichol, 2016) and for object-driven generation under Dreambooth dataset (Ruiz et al., 2023) under different fine-tuning methods. $\mbox{A-ESR}_{0.6}$ and $\mbox{A-ESR}_{0.7}$ refer to the Average Extraction Success Rate, with the threshold $\tau$ for successful extraction set at 0.6 and 0.7, respectively. The experimental results demonstrate that FineXtract exhibits stronger performance under all scenarios and metrics than baselines.
@@
4.3 Clustering Generated Images
Specifically, inspired by Carlini et al. (2023), we compute the similarity between each pair of generated images and construct a graph where each image is represented as a vertex. We connect two vertices when the similarity between the corresponding images exceeds a threshold $\phi$, i.e., if $\text{sim}(x_{i},x_{j})\geq\phi$, we connect vertices $i$ and $j$. By default, we use SSCD (Pizzi et al., 2022) to measure similarity, in line with previous work (Somepalli et al., 2023a, b). Instead of using a fixed threshold (Carlini et al., 2023), we gradually increase the threshold $\phi$ until the number of cliques, each denoted by $A^{(k)}$, within the graph identical to the proposed number of training images $N_{0}$. This approach helps us identify the generated image subset (i.e., the clique $A^{(k)}$ ) corresponding to each training image ($X_{0}^{(k)})$. Next, we identify the central image $\hat{x}^{(k)}$ for each clique $A^{(k)}$. The central image is defined as the one that maximizes the average similarity with the other images in the clique: $\hat{x}^{(k)}=\arg\max\limits_{x}\frac{1}{|A^{(k)}|}\sum_{x_{q}\in A^{(k)}}\text{sim}(x,x_{q}).$ The final extracted image set is represented as $\widehat{X}=\{\hat{x}^{(0)},\hat{x}^{(1)},\dots,\hat{x}^{(N_{0})}\}$.
5 Experiments
In this section, we apply our proposed method, FineXtract, to extract training data under various few-shot fine-tuning techniques across different types of DMs. We conduct experiments on two common scenarios for few-shot fine-tuning: style-driven and object-driven generation. For style-driven generation, which focuses on capturing the key style of a set of images, we randomly select 20 artists, each with 10 images, from the WikiArt dataset (Nichol, 2016). For object-driven generation, which emphasizes the details of a given object, we experiment on 30 objects from the Dreambooth dataset (Ruiz et al., 2023), each consisting of 4-6 images. This setup aligns with the recommended number of training samples in the aforementioned fine-tuning methods (Ruiz et al., 2023; Hu et al., 2021). We experiment with two most widely-used few-shot fine-tuning techniques: DreamBooth (Ruiz et al., 2023), and LoRA (Hu et al., 2021). More details for the fine-tuning setting are available in Appendix Sec. D.
@@ 5 Experiments
The default model used for training is Stable Diffusion (SD) V1.4222https://huggingface.co/CompVis/stable-diffusion-v1-4. Additionally, we demonstrate the adaptability of our method to various types and versions of DMs, larger training datasets, and different numbers of generated images (refer to Sec.  5.2 for more details).
@@ 5.1 Comparison
Previous extraction methods primarily focus on the generation capabilities of text-to-image DMs, employing either direct text-to-image generation or classifier-free guidance (CFG) (Carlini et al., 2023; Somepalli et al., 2023a, b). To better demonstrate the effectiveness of our framework, we compare FineXtract with Direct Text-to-Image and CFG, both combined with the clustering algorithm proposed in Section 4.3. For both CFG and FineXtract, we set the guidance scale $w^{\prime}$ to 3.0 under DreamBooth fine-tuning. Under LoRA fine-tuning, $w^{\prime}$ are set to 3.0 for CFG and 5.0 for FineXtract. These hyperparameters are found to perform well (see Sec. 5.3 for details). All methods use the same number of generation iterations, $N$, set to $50\times N_{0}$, and the number of extracted images set to $N_{0}$ to ensure a fair comparison. The results, shown in Table 1, demonstrate a significant advantage of FineXtract over previous methods, with an improvement of approximately 0.02 to 0.05 in AS and a doubling of the A-ESR in most cases.
Our method, as detailed in Sec. 4.1 and 4.2, employs guidance between two distinct models at each generation step without slowing down the generation speed compared to traditional CFG (which also requires guidance by forwarding the main UNet in DMs twice). However, GPU memory costs increase due to the need to load both the pretrained and fine-tuned models. In Tab. 2, we present a demo experiment comparing the computational costs of FineXtract and CFG, using SD (v1.4) fine-tuned with DreamBooth and LoRA on the WikiArt dataset. The batch size is fixed at 5, and all experiments are conducted on a single A100 GPU.
Table 2: Generation GPU memory costs and time costs for per-image generation using FineXtract and CFG.
[t]
Metrics and Settings DreamBooth LoRA
Memory Costs (MB) Time Costs (s) Memory Costs (MB) Time Costs (s)
FineXtract 15812 3.4 15894 3.6
CFG 12096 3.4 12226 3.6
@@ 5.2 Generalization
(a) Different $N_{0}$
(b) Different NN
Figure 4: Experiment on generalization ability of FineXtract across different number of generated images and training images. We experiment on 4 classes of WikiArt fine-tuning on SD (V1.4) with DreamBooth. FineXtract consistently outperforms baseline under different NN and $N_{0}$.
@@
Figure 5: Visualization of generated images using FineXtract under different $w^{\prime}$ with fixed $x_{T}$.
Number of Training Images $N_{0}$. As the number of training images increases, the learned concept during fine-tuning becomes more intricate, thereby enhancing the difficulty of extracting training images. To thoroughly examine how this influences performance, we conduct experiments with varying numbers of training images in 5 classes in WikiArt, the results of which are presented in Fig. 4(a), where we can observe a performance drops when the number of the training images is large.
Different Numbers of Generated Images $N$. As previously mentioned, the clustering algorithm allows attackers to leverage more generation iterations to improve extraction accuracy. We test our method with different numbers of generated images $N$, ranging from $N_{0}$ to $50\times N_{0}$. As shown in Fig. 4(b), increasing $N$ significantly improves performance. However, the time complexity of finding maximal cliques can grow exponentially with the number of nodes (Tomita et al., 2006). Thus, further increasing $N$ to larger than $50\times N_{0}$ makes it considerably more difficult for the clustering algorithm to converge.
5.3 Ablation Study
(a) Ablation on $w^{\prime}$
(b) Ablation on $k$
Figure 6: Ablation Study on hyper-parameters $w^{\prime}$ and $k$. We experiment on 4 classes of WikiArt dataset fine-tuning on SD (V1.4) with DreamBooth.
In this section, we experiment with hyperparameters in Eq. 13, including the guidance scale $w^{\prime}$ and the correction term scale $k$. We experiment on 4 classes in WikiArt fine-tuning with DreamBooth. Results under LoRA are shown in Appendix Sec. C.
Guidance Scale $w^{\prime}$. The guidance scale $w^{\prime}$ is the most critical hyperparameter influencing extraction efficiency. If $w^{\prime}$ is too low, the guidance provided by fine-tuning methods is weakened. Conversely, if $w^{\prime}$ is too high, it often causes generation failures, resulting in unrealistic outputs (see visual examples in Appendix Fig. 5). As shown in Fig. 6(a), $w^{\prime}=3.0$ works well for both CFG and FineXtract when DMs are fine-tuned using DreamBooth.
Correction Term Scale $k$. In Eq. 13, we derive a correction term $k\epsilon_{\theta}(x_{t},t)$ to address the inconsistency between CFG and model guidance. Although Eq. 13 indicates that $k$ should not be less than 0, our experiments suggest that setting k=0.02k=-0.02 typically yields the best results, as illustrated in Fig. 6(b). This may be due to the complex interaction between $w^{\prime}$ and $k$, where $k$ can only be guaranteed to be greater than 0 if the true value of $w^{\prime}$ can be identified, which is often impractical in real-world scenarios.